# Information Sifting Funnel: Privacy-preserving Collaborative Inference Against Model Inversion Attacks

### v1

@@ 4.3. Experimental Demonstration of the Criterion and Guidance for Defense Design
Where $p(x,z)$ represents the joint distribution, existing defense methods approximate it, but limited batch sizes lead to inaccuracies and variations in mutual information estimates. For high-dimensional data (e.g., images), accurate estimation requires many samples.
To avoid direct distribution estimation, the Donsker-Varadhan representation reformulates mutual information as an optimization problem:
@@
Where $T$ is a learnable function, this formulation enables mutual information to be estimated by optimizing $T$ to obtain a lower bound.
@@
As shown in Figure 2, we trained target models on the CIFAR-10 (Krizhevsky et al., 2009) and used Equation (11) to train MINE for the corresponding blocks. Figures 2(a, b) show that increasing network depth reduces both mutual information and effective information, leading to a decline in attack performance. Figures 2(c, d) illustrate that while skip connections within two layers do not reduce effective information, they slightly lower mutual information, moderately affecting attack performance.
@@
Previous research (Hao et al., 2025) identified vulnerabilities caused by skip connections in centralized MIA. Our study further highlights their impact on MIA effectiveness in collaborative inference. Transmission between layers alters feature channels and dimensions, requiring a shortcut layer, which leads to a more pronounced decline in mutual information compared to direct stacking within layers. This, in turn, degrades MIA performance and aligns with our criterion.
Notably, while mutual information at ResNet’s fourth layer is higher than that at CNN’s second layer, MLE-based MIA does not perform better. This is because these methods rely heavily on gradient descent, making them sensitive to suboptimal optimizers, hyper-parameter settings, and deeper network structures, which hinder gradient calculations for optimizing $x$. The lack of flexible and efficient methods for MLE-based MIA remains a gap in current research, often leading to their exclusion in defense evaluations—a topic warranting further exploration (Ding et al., 2024).
In our experiments, we still evaluate MLE-based MIA against various defenses and provide evidence demonstrating the effectiveness of our method in defending against such attacks.
In conclusion, through theoretical analysis and experimental validation, we established $D_{mia}$ and determined that effective defense strategies must impact mutual information, entropy, and $\delta(z)$. The following considerations should guide defense design:
@@ 5. SiftFunnel: Information Sifting Funnel
Building on the theoretical analysis, criterion, and experiments in Section 4, this section introduces SiftFunnel and its adherence to $D_{mia}$ in edge model and loss function design.
We first outline the core principles and objectives of SiftFunnel. Then, we detail the edge model’s structure, key modules, and role in the forward propagation of information. Finally, we explain the loss function’s impact on $H(x\mid z)$, $H(z)$ and $\delta(z)$, highlighting its ability to extract useful information while minimizing redundancy.
@@ 5.1. Overview of SiftFunnel
Guided by $D_{{mia}}$, SiftFunnel aims to balance usability and privacy in edge models by increasing the uncertainty of inferring $x$ from $z$ while reducing $z$’s information content and separability. To this end, as shown in Figure 3, we introduce two key improvements: First, the edge model is enhanced by modifying the final hidden layer into a funnel-shaped structure and incorporating attention mechanism modules between layers. Second, the loss function is optimized to ensure that model training balances usability with privacy protection.
Specifically, in the edge model design, the final layer adopts a funnel-shaped structure, reducing feature channels while preserving spatial dimensions. This reduces computation, lightens the edge device load, and limits redundant information. Attention mechanisms and multi-receptive field convolutions enhance task-relevant information, while a feature channel supplementation module in the cloud model ensures task accuracy. With the improved edge model, designing an effective loss function is essential. To reduce redundancy and mutual information between $z$ and $x$, we use Distance Correlation (Li et al., 2012) for nonlinear correlations and the Pearson Coefficient (Cohen et al., 2009) with $l_{1}$-norm for linear correlations. Additionally, label smoothing and KL divergence (Donsker and Varadhan, 1983) ensure model accuracy, enhancing usability while reducing the separability of $z$.
@@ 5.2. The Design of Shallow Edge Model
Information Funnel. Deepening the neural network enhances feature extraction and reduces $I(x,z)$, as confirmed in the Section 4. Previous work (Ho et al., 2024) also demonstrates through Fisher information (Rissanen, 1996) that deeper networks improve feature relevance to tasks. Inspired by this, we redesign the final layer of the edge model with an Information Funnel module. This module reduces output channels to one-fourth in the first layer and further in the second layer, with channel selection based on the spatial size of feature maps. If the spatial size is reduced significantly, channels can exceed 10 to maintain accuracy. If unchanged, channels can be set to 2 to limit redundancy while preserving accuracy. Adjusting channels based on spatial size is necessary because spatial information loss is irreversible, whereas channel information loss can be compensated in the cloud model. While this compensation increases $I(z,y)$, it does not reverse the reduction in $I(x,z)$, ensuring attack difficulty remains high, as demonstrated in Section 6.3. So, we add the channel compensation module in the cloud model to guarantee usability. Notably, assumption about extracting features from non-independent transmission modules is extremely strong and idealized, requiring the adversary to have full white-box knowledge of the cloud model as if they were the developer.
[…]
3
@@
Attention Block. To enhance the output of task-relevant information from the Information Funnel, attention mechanism modules are added both before and after it. The first module combines an SE block (Hu et al., 2018) and Triplet Attention (Misra et al., 2021). The SE block evaluates the contribution of information within each channel, suppressing or nullifying channels with low task relevance. Triplet Attention reduces spatial information while maintaining smooth and continuous outputs, preserving fine-grained features suitable for further processing. The second attention mechanism integrates SE and CBAM modules (Woo et al., 2018). Compared to Triplet Attention, CBAM is a more lightweight module that sequentially calculates channel and spatial weights to enhance or suppress input features, improving network performance. Since CBAM’s channel and spatial attention are independently designed, certain regions are suppressed redundantly, making the output features more sparse. This is ideal for final output processing, effectively increasing $H(x|z)$ while suppressing $H(z)$ and $\delta(z)$.
@@ 5.3. The Design of Loss Function
To further enhance model usability and data privacy, a appropriately designed loss function is essential to guide the training process.
To reasonably increase $H(x\mid z)$ and suppress $H(z)$ within the edge model structure—thereby increasing the uncertainty of inferring $x$ from $z$—loss constraints must be applied to $x$ and $z$. The key challenge is that, without intervention, neural networks may retain irrelevant input information in their output during forward propagation. This information, which does not significantly impact decision-making, reduces the difficulty of attacks by maintaining $I(x,z)$.
Some defense approaches utilize Constrained Covariance (COCO) (Gretton et al., 2005b) to estimate correlation, focusing primarily on the linear relationship between $x$ and $z$. However, since neural networks predominantly perform nonlinear operations, studies have shown that this constraint is not only ineffective but may also lead to reduced accuracy (Peng et al., 2022). Alternatively, some approaches employ the Hilbert-Schmidt Independency Criterion (HSIC) (Gretton et al., 2005a) to estimate correlation. However, kernel-based calculations in HSIC rely heavily on the appropriate selection of $\sigma$. Improper $\sigma$ value can introduce bias in nonlinear correlation estimation, limiting the generalizability of this method and reducing its effectiveness across various scenarios.
Nonlinear correlation constraint. We observe that shallow neural network features retain significant similarity to the input and are derived through nonlinear operations. To address this, we use distance correlation to measure the nonlinear relationship between $x$ and $z$, incorporating it as a loss constraint. This helps remove redundant input information from the output, effectively increasing $H(x\mid z)$. The calculation of distance correlation is as follows:
@@
Where $\mathrm{d}\operatorname{Cov}(x,z)$ represents the distance covariance between $x$ and $z$, and the denominator is the product of the distance variances for the two datasets.
@@
The distance variances is defined as $\mathrm{d}\operatorname{Var}(x)=\frac{1}{\mathrm{n}^{2}}\sum_{\mathrm{i},\mathrm{j}}A_{ij}^{c^{2}}$, which normalizes the distance covariance. The value range of distance correlation is $[0,1]$, where $1$ indicates perfect correlation. Since $x$ and $z$ exhibit high distance similarity and are derived through nonlinear operations, distance correlation is well-suited as a constraint term. Moreover, as a distribution-free measure, it does not rely on specific data distributions, making it effective for non-Gaussian distributions, nonlinear relationships, and complex data types such as images and text. For non-stationary and multimodal data, distance correlation provides stable dependency measurements, allowing the constraint to flexibly adapt to different models and datasets.
Linear correlation constraint. In addition to constraining the loss between $x$ and $z$, it is necessary to design loss constraints specifically for the $z$. The primary issue lies in the $z$ containing repetitive or similar information, which is redundant for decision-making and provide additional input for Gen-based MIAs, strengthening local constraints in the input space and improving reconstruction accuracy.
To address this, we propose using the Pearson correlation coefficient as a loss constraint to reduce the linear correlation within features $z$, effectively eliminating redundant information. The calculation is as follows:
@@
To compute the scalar $\mathcal{L}_{\text{Pearson }}$, the squared correlations of all channel pairs are averaged, keeping the value range within $[0,1]$. This avoids negative values, which could lead to suboptimal updates during gradient descent. Minimizing the $\mathcal{L}_{\text{Pearson }}$ in the presence of negative values might push the correlation toward complete negative correlation, introducing unnecessary feature redundancy or unsuitable learning patterns, ultimately preventing convergence. Therefore, the goal is to minimize the coefficient toward zero, as lower values indicate reduced linear correlation. Using the squared correlations and their average ensures a smoother and more stable gradient descent process.
Additionally, to prevent extreme values in the feature output weights during the optimization of the above constraints, we include the $l_{1}$-norm as a loss term. This term is added specifically to suppress weights, but excessive emphasis on it could interfere with normal model training. Therefore, the coefficient $\tau$ for this loss term is set below 1e-2 to balance its influence.
Label smoothing. Finally, to ensure model accuracy, we draw inspiration from (Struppek et al., 2023) and apply label smoothing to refine the target loss, as shown in the formula (14). Additionally, KL divergence is used instead of cross-entropy (CE) as the primary loss function for optimization.
@@
This smoothing mechanism ensures that the loss value computed by cross-entropy emphasizes enhancing inter-class feature distinctions, while KL divergence as a loss function supports smoother and more generalized feature extraction between classes (Struppek et al., 2023). Within-class feature extraction focuses on capturing universal characteristics, ensuring the model maintains a certain level of generalization and robustness for the target task. Moreover, this optimization reduces maximum feature separability, increasing the difficulty of attacks, as demonstrated in Section 6.
In summary, the proposed loss function first applies nonlinear correlation constraints between input $x$ and features $z$ to increase $H(x\mid z)$ and suppress $H(z)$. Second, it imposes linear correlation and extreme value constraints on $z$ to further enhance $H(x\mid z)$ and suppress $\delta(z)$. Finally, label smoothing is applied to the target loss to maintain accuracy while slightly suppressing $H(z)$. The final formulation of this optimization is as follows, with specific parameter settings referenced in Section 6.
@@ 6. Experiments
In this section, we systematically evaluate the resistance to MIA and compare our method with existing approaches. The results are analyzed from three perspectives: the impact of defense methods on $I(x\mid z)$ and $\delta(z)$, the influence of datasets and model architectures, and ablation studies. The specific experimental setup and evaluation metrics are shown below.
@@ 6.1. Experimental Setup
  • CIFAR-10 (Krizhevsky et al., 2009). The CIFAR-10 dataset consists of 60,000 RGB images categorized into ten classes (e.g., airplane, ship, etc.) with an original resolution of $32\times 32$. To make the results clearer and more intuitive, we used an upscaled resolution of $64\times 64$.
@@
  • CelebA (Liu et al., 2015). CelebA is a dataset with 202,599 images of 10,177 celebrities from the Internet. We used the same crop as (Yang et al., 2019) to remove the background of images in this dataset other than faces to reduce the impact on the experiment. In order to eliminate individual overlap, we removed a total of 6,878 images of 296 individuals from CelebA and similarly resized the images to $64\times 64$.
  • ChestX-ray (Sait et al., 2020). This dataset is a curated collection of COVID-19 chest X-ray images compiled from 15 publicly available datasets. It contains 1,281 COVID-19 X-rays, 3,270 normal X-rays, 1,656 viral pneumonia X-rays, and 3,001 bacterial pneumonia X-rays. For our analysis, we converted the images to gray scale with a resolution of $128\times 128$.
@@
Classifier Attack Model
@@
2) Attack Method. The selection of attack methods has been discussed in Section 3.2. Specifically, for MLE-based MIA, this paper has adjusted to the hyperparameter values that yield the best results. For Gen-based MIA, the attack model in this paper is designed as the inverse architecture of the edge model. Moreover, under the gray-box scenario assumption, the attack model can enhance the expressive power of the reverse mapping as the edge model architecture changes. The attack model is also trained using the Adam optimizer with a learning rate of 2e-4, a $\beta_{1}$ of 0.5, and a ReduceLROnPlateaud scheduler with a factor of 0.5 and patience of 20. The training conditions for the attack models in this paper have been done with strong assumptions and only the attack models where the test Mean Square Error (MSE) reaches the minimum value are saved to highlight the effectiveness of the defense efforts.
3) Target Edge Model and Implementation Details. We employ the same comprehensive model architecture as detailed in (Yang et al., 2019), comprising four Convolutional Neural Network (CNN) blocks succeeded by two fully connected layers. This model is trained using the Adam optimizer with a learning rate of 2e-4, a $\beta_{1}$ parameter of 0.5, and a ReduceLROnPlateaud scheduler with a reduction factor of 0.5 and a patience setting of 25. Each CNN block is composed of a convolutional layer, a batch normalization layer, a max-pooling layer, and a ReLU activation function. The target edge model encompasses the initial two CNN blocks. For the training of the edge model protected by SiftFunnel, we have specified the parameters $\alpha$ to 3.5, $\epsilon$ to 0.35, $\beta$ to 0.8, and $\gamma$ to 0.6 in Equation (15). In addition to this, to argue the impact of the model architecture for the defense, we use different CNNs as follows: (1) VGG16 adapted from (Simonyan, 2014); (2) ResNet-18 adapted from (He et al., 2016), the learning rates are 3e-4. The division positions of resnet and vgg are not fixed. For detailed accuracy and how to divide them, see Section 6.3 for analysis. All experiments were performed on two RTX 4090 GPUs and an Intel Core i9-14900KF ×$\times 32$ CPU.
4) Comparison of Defense Methods. We will compare three existing categories of defense methods. For Perturbation-based defenses, we select Adding noise & Dropping features (He et al., 2020), DPSGD (Abadi et al., 2016), and negative label smoothing (NLS) (Struppek et al., 2023). For IB-based defenses, we choose MID (Wang et al., 2021), BiDO (with HSIC) (Peng et al., 2022), VIB (Wang et al., 2024), and InfoSCISSORS (Duan et al., 2023). For NND & IB-based defenses, we opt for AE-based protection (Azizian and Bajić, 2024). Our method falls into the NND & IB-based defense category, which also requires analyzing the defensive effect in a gray-box scenario. Furthermore, all comparison methods have been selected and set with appropriate hyper-parameters for comparison, with detailed settings found in Section 6.3 analysis. Lastly, all methods are trained without adversarial training, focusing solely on the effects of the original defense methods.
6.2. Evaluation Metrics
We have selected 7 evaluation metrics to assess usability, privacy, and deployability. For usability assessment, we employ Test Accuracy (Test ACC), which evaluates the final performance of the model protected by defense techniques on the test dataset. Regarding privacy assessment, we initially utilize MSE, Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM) to evaluate the effectiveness of MIA on reconstructing the input data of the protected model (Yin et al., 2023). Here, a higher MSE and lower PSNR and SSIM (ranging from 0 to 1) indicate better defense effectiveness. Compared to the former two, SSIM focuses more on the perceptual quality of the image, encompassing luminance, contrast, and structural information. In addition to these, we also employ MINE (Belghazi et al., 2018), as mentioned in Section  4.3, to assess Mutual Information (MI) (ranging from 0 to 1) and the Effective Information Mean $\delta(z)$, which is the mean of the non-zero information output by the edge model on the test dataset. The specific training methods and model architecture for MINE are referenced from xxx. For deployability, we consider the Edge model’s parameter count $|\theta_{\text{edge}}|$ and Inference Latency. The model parameter count directly determines memory or GPU memory occupation, while inference latency is tested on both CPU and GPU.
6.3. Experimental Results
@@
MLE-based Attack (He et al., 2016) Gen-based Attack (He et al., 2020)
@@
Table 3. Quantitative evaluation of the resistance to MIA across different architectures on CIFAR-10. The edge model divisions for CNN and ResNet-18 follow the configurations in Tables 2 and 4, while for VGG16, the edge model is set at the first ReLU layer after a max-pooling operation.

### v4

@@ 4.3. Demonstration and Guidance for Defense
Where $p(x,z)$ represents the joint distribution, existing defense methods approximate it, but limited batch sizes lead to inaccuracies and variations in estimates. For high-dimensional data (e.g., images), accurate estimation requires many samples.
To avoid direct distribution estimation, the DV representation reformulates mutual information as an optimization problem:
@@
Where $T$ is a learnable function, this formulation enables $I(X;Z)$ to be estimated by optimizing $T$ to obtain a lower bound.
@@
As shown in Figure 2, we trained the target models on the CIFAR-10 (Krizhevsky et al., 2009) and used Equation (11) to train MINE for the corresponding blocks. Figures 2(a, b) show that increasing network depth reduces both mutual information and effective information, leading to a decrease in attack performance. Figures 2(c, d) illustrate that while skip connections within two layers do not reduce $\delta(z)$, they slightly lower $I(x,z)$, moderately affecting attack performance. Previous research (Hao et al., 2025) identified vulnerabilities caused by skip connections in centralized MIA. Our study further highlights their impact on MIA effectiveness in CI.
@@
However, Transmission between blocks in ResNet alters feature channels and dimensions, requiring a shortcut layer, leading to a more pronounced decline in mutual information compared to direct stacking within layers. This, in turn, degrades MIA performance and aligns with our criterion.
[…]
3
Notably, while mutual information at ResNet’s fourth layer is higher than that at CNN’s second layer, MLE-based MIA does not perform better. This stems from the fact that these methods rely heavily on gradient descent, making them sensitive to suboptimal optimizers, hyper-parameter settings, and deeper network structures, which hinder gradient calculations for optimizing $x$. The lack of flexible and efficient methods for MLE-based MIA remains a gap in current research, often leading to their exclusion in defense evaluations—a topic warranting further exploration (Ding et al., 2024). In our experiment and Section 6, we still evaluate MLE-based MIA against various defenses and provide evidence demonstrating the effectiveness of our method in defending against such attacks.
In conclusion, through theoretical analysis and experimental validation, we proposed $D_{mia}$ and determined that effective defense strategies must impact mutual information, entropy, and $\delta(z)$. The following considerations can guide defense design:
@@ 5.1. Overview of SiftFunnel
Guided by $D_{{mia}}$, SiftFunnel balances usability and privacy by increasing uncertainty in inferring $x$ from $z$ while limiting redundant information and separability. As shown in Figure 3, SiftFunnel introduces two key improvements: First, the edge model modifies the final hidden layer into a funnel-shaped structure and incorporates attention modules. Second, the loss function is optimized to ensure that model training balances usability with privacy protection.
Specifically, in the edge model design, the final layer adopts a funnel-shaped structure, reducing feature channels while preserving spatial dimensions. This reduces computation, lightens the edge device load, and limits redundant information. Attention mechanisms and multi-receptive field convolutions enhance task-relevant information, while a cloud-based upsampling module preserves accuracy. To reduce redundancy, $I(x,z)$ and increase the uncertainty of $z$ to $x$, we use Distance Correlation (Li et al., 2012) for nonlinear correlations and the Pearson Coefficient (Cohen et al., 2009) with $l_{1}$-norm for linear correlations. Additionally, Label Smoothing (LS) and Kullback-Leibler (KL) divergence (Donsker and Varadhan, 1983) are used during training to improve model usability while reducing the separability of representations in $z$.
@@ 5.2. The Design of Shallow Edge Model
Information Funnel. Deepening the neural network enhances feature extraction and reduces $I(x,z)$, as confirmed in Section 4. Previous work (Ho et al., 2024) shows through Fisher information (Rissanen, 1996) that deeper networks improve task relevance and resist MIA. Inspired by this, we redesign the final layer of the edge model with an Information Funnel module. This module reduces output channels to one-fourth in the first layer and further in the second layer, with channel selection based on the spatial size of feature maps. If the spatial size is reduced significantly, channels can exceed 10 to maintain accuracy. If unchanged, channels can be set to 2 to limit redundancy while preserving accuracy. Adjusting channels based on spatial size is necessary because spatial information loss is irreversible, whereas channel information loss can be compensated in the cloud model. This compensation increases $I(z,y)$ without reversing the reduction in $I(x,z)$, maintaining high attack difficulty as shown in Section 6.3. So, we add the channel compensation module in the cloud model to guarantee usability. Importantly, assumptions about extracting features from non-independent transmission modules are idealized, requiring adversaries to have full white-box knowledge of the cloud model as if they were the developer.
@@
Attention Block. To enhance the output of task-relevant information from the Information Funnel, attention modules are incorporated both before and after it. The first block combines an SE block (Hu et al., 2018) and Triplet Attention (Misra et al., 2021). The SE block evaluates the contribution of information within each channel, suppressing channels with low task relevance. Triplet Attention reduces spatial information while maintaining smooth, continuous outputs, preserving fine-grained features for further processing. The second block integrates SE and CBAM modules (Woo et al., 2018). Compared to Triplet Attention, CBAM is a lightweight module that sequentially calculates channel and spatial weights to enhance or suppress input features. Since CBAM’s channel and spatial attention are independently designed, certain weights are suppressed redundantly, resulting in sparser outputs. This sparsity is beneficial for output processing, effectively increasing $H(x\mid z)$ while reducing $H(z)$ and $\delta(z)$.
@@ 5.3. The Design of Loss Function
To further enhance model usability and data privacy, appropriately designed loss constraints must be applied to $x$ and $z$ to reasonably increase $H(x\mid z)$ and suppress $H(z)$ within the edge model structure. Without intervention, neural networks may retain irrelevant input information in their output during forward propagation. This information, which does not significantly impact decision-making, reduces the difficulty of MIAs by maintaining the mutual information and entropy. The key challenge is how to achieve this effectively and flexibly.
Some defense approaches utilize Constrained Covariance (COCO) (Gretton and Herbrich, 2005) to estimate correlation, focusing primarily on the linear relationship between $x$ and $z$. However, since neural networks predominantly perform nonlinear operations, studies have shown that this constraint is not only ineffective but may also lead to reduced usability (Peng et al., 2022). Alternatively, some approaches employ the Hilbert-Schmidt Independency Criterion (HSIC) (Gretton et al., 2005) to estimate nonlinear correlation. However, kernel-based calculations in HSIC rely heavily on the appropriate selection of $\sigma$. Improper $\sigma$ value can introduce bias in correlation estimation, limiting the generalizability of this method and reducing its effectiveness across various scenarios.
Nonlinear correlation constraint. We observe that shallow neural network features retain significant similarity to the input and are derived through nonlinear operations. To address this, we use distance correlation to measure the non-linear relationship between $x$ and $z$, incorporating it as a loss constraint. This helps remove redundant input information from the output and effectively increase $H(x\mid z)$. The calculation of distance correlation is as follows:
@@
Where $\mathrm{d}\operatorname{Cov}(x,z)$ represents the distance covariance between $x$ and $z$, and the denominator is the product of the distance variances for the two samples.
@@
The distance variance is defined as $\mathrm{d}\operatorname{Var}(x)=\frac{1}{\mathrm{n}^{2}}\sum_{\mathrm{i},\mathrm{j}}A_{ij}^{c^{2}}$, which normalizes the distance covariance. The value range of distance correlation is $[0,1]$, where $1$ indicates perfect correlation. Since $x$ and $z$ exhibit high distance similarity and are derived through nonlinear operations, distance correlation is well-suited as a constraint term. Moreover, as a distribution-free measure, it does not rely on specific data distributions, making it effective for non-Gaussian distributions, nonlinear relationships, and complex data types such as images and text. For non-stationary and multimodal data, distance correlation provides stable dependency measurements, allowing the constraint to flexibly adapt to different models and data.
Linear correlation constraint. In addition to constraining the correlation between $x$ and $z$, it is necessary to design loss constraints specifically for the $z$. The primary issue lies in the $z$ containing repetitive or similar information, which is redundant for decision-making and provides additional input for Gen-based MIAs, strengthening local constraints in the input space and improving reconstruction accuracy.
To address this, we propose using the Pearson correlation coefficient as a loss constraint to reduce the linear correlation between $x$ and $z$, effectively eliminating redundant information and reducing $\delta(z)$. The calculation is as follows:
@@
To compute the scalar $\mathcal{L}_{\text{Pearson }}$, the squared correlations of all channel pairs are averaged, keeping the value range within $[0,1]$. This avoids negative values, which could lead to suboptimal updates during gradient descent. with negative values might drive correlations toward complete negative correlation, introducing unnecessary redundancy or unsuitable learning patterns and hindering convergence. Therefore, the goal is to minimize the coefficient toward zero, as lower values indicate reduced linear correlation. Using the squared correlations and their average ensures a smoother and more stable gradient descent process.
Additionally, to prevent extreme values in the feature output weights during the optimization of the above constraints, we include the $l_{1}$-norm as a loss term. This term is added specifically to suppress weights, but excessive emphasis could interfere with normal model training. Therefore, the coefficient $\tau$ for this loss term is set below 1e-2 to balance its influence.
Label smoothing. Finally, to ensure model usability, we draw inspiration from (Struppek et al., 2023) and apply LS to refine the target loss, as shown in formula (14). Additionally, KL divergence is used instead of cross-entropy (CE) as the primary loss function for optimization.
@@
CE enhances inter-class feature distinctions, while KL divergence supports smoother and more generalized feature extraction between classes (Struppek et al., 2023). Within-class feature extraction captures universal characteristics, ensuring generalization and robustness for the target task. This optimization reduces feature separability, increasing the difficulty of attack, as demonstrated in Section 6.3.
In summary, the proposed loss function first applies nonlinear correlation constraints to increase $H(x|z)$ and suppress $H(z)$. Next, it applies linear correlation and extreme value constraints on $z$ to further enhance $H(x\mid z)$ and suppress $\delta(z)$. Finally, LS and KL divergence maintain usability while slightly suppressing $H(z)$. The final formulation of this optimization is as follows, with specific parameter settings. The final formulation of this optimization is as follows, with specific parameter settings referenced in Section 6.1. SiftFunnel and MINE are available at https://github.com/SiftFunnel/SiftFunnel.
@@ 6. Experiments
In this section, we systematically evaluate and compare the ability of SiftFunnel and existing approaches to resist MIA. The specific experimental setup and evaluation metrics are shown below.
@@ 6.1. Experimental Setup
  • CIFAR-10 (Krizhevsky et al., 2009). The CIFAR-10 dataset consists of 60,000 RGB images categorized into ten classes with an original resolution of $32\times 32$. To make the results clearer and more intuitive, we used an unscaled resolution of $64\times 64$.
@@
  • CelebA (Liu et al., 2015). CelebA is a dataset with 202,599 images of 10,177 celebrities. We used the same crop as (Yang et al., 2019) to remove the background of images in this dataset other than faces to reduce the impact on the experiment. To eliminate individual overlap, we removed a total of 6,878 images of 296 individuals and similarly resized the images to $64\times 64$.
  • ChestX-ray (Sait et al., 2020). This dataset is a curated collection of COVID-19 chest X-ray images compiled from 15 publicly available datasets. It contains 1,281 COVID-19 X-rays, 3,270 normal X-rays, 1,656 viral pneumonia X-rays, and 3,001 bacterial pneumonia X-rays with a resolution of $128\times 128$.
@@
Target model Attack Model
@@
MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
@@
2) Attack Method. The selection of attack methods has been discussed in Section 3.2. Specifically, for MLE-based MIA, this paper has adjusted to the hyper-parameter values that yield the best results. For Gen-based MIA, the attack model in this paper is designed as the inverse architecture of the edge model. Moreover, under the gray-box scenario assumption, the attack model can enhance the expressive power of the reverse mapping as the edge model architecture changes. The attack model is trained using the Adam optimizer with a learning rate of 2e-4, a $\beta_{1}$ of 0.5, and a ReduceLROnPlateaud scheduler with a factor of 0.5 and patience of 20. The attack model training is based on strong assumptions, as shown in Table 1, and only the models with the minimum test MSE are saved to highlight the defense’s effectiveness.
3) Target Edge Model and Implementation Details. We employ the same comprehensive model architecture as detailed in (Yang et al., 2019), comprising four CNN blocks succeeded by two fully connected layers. This model is trained using the Adam optimizer with a learning rate of 2e-4, a $\beta_{1}$ parameter of 0.5, and a ReduceLROnPlateaud scheduler with a reduction factor of 0.5 and patience of 25. Each CNN block is composed of a convolutional layer, a batch normalization layer, a max-pooling layer, and a ReLU activation function. The target edge model encompasses the initial two CNN blocks. For the training of the edge model protected by SiftFunnel, we have specified the parameters $\lambda_{1}$ to 3.5, $\alpha$ to 0.35, $\lambda_{2}$ to 0.8, and $\lambda_{3}$ to 0.6 in Equation (15). In addition, to argue the impact of the model architecture for the defense, we use different CNNs as follows: (1) VGG16 adapted from (Simonyan, 2014); (2) ResNet-18 adapted from (He et al., 2016), the learning rates are 3e-4. The division positions of ResNet and VGG16 are not fixed. For detailed accuracy and how to divide them, see Section 6.3 for analysis. All experiments were performed on two RTX 4090 GPUs and an Intel Core i9-14900KF ×$\times 32$ CPU.
4) Comparison of Defense Methods. We aim to compare three existing categories of defense methods. For Perturbation-based defenses, we select Adding noise & Dropping features (He et al., 2020), DPSGD (Abadi et al., 2016), and negative label smoothing (NLS) (Struppek et al., 2023). For IB-based defenses, we choose MID (Wang et al., 2021), BiDO (with HSIC) (Peng et al., 2022), VIB (Wang et al., 2024), and InfoSCISSORS (Duan et al., 2023). For NND & IB-based defenses, we opt for AE-based protection (Azizian and Bajić, 2024). Our method falls into the NND & IB-based defense category, which also requires analyzing the defensive effect in a gray-box scenario. Furthermore, all comparison methods have been selected and set with appropriate hyper-parameters for comparison, with detailed settings found in Section 6.3. Lastly, all methods are trained without adversarial training, focusing solely on the effects of the original defense methods.
6.2. Evaluation Metrics
We selected 7 evaluation metrics to assess the aspects of usability, privacy, and deployability. For usability assessment, we employ Test Accuracy (Test ACC), which evaluates the final performance of the model protected by defense techniques on the test dataset. Regarding privacy assessment, we initially utilize MSE, Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM) to evaluate the effectiveness of MIA on reconstructing the input data of the protected model (Yin et al., 2023). Here, a higher MSE and lower PSNR and SSIM (ranging from 0 to 1) indicate better defense effectiveness. In addition, we employ MINE (Belghazi et al., 2018), as mentioned in Section  4.3, to assess Mutual Information (MI) (ranging from 0 to 1) and the Effective Information Mean $\delta(z)$, which is the mean of the non-zero output by the edge model on the test dataset. The specific training methods and model architecture for MINE are referenced from https://github.com/SiftFunnel/SiftFunnel. For deployability, we consider the Edge model’s parameter count $|\theta_{\text{edge}}|$ and Inference Latency. The model parameter count directly determines memory or GPU memory occupation, while inference latency is tested on both CPU and GPU.